OpenStudy (anonymous):

What is the value of y when x = 3? y = 12x x y 0 1 2 3 A. 24 B. 36 C. 48 D. 60

OpenStudy (anonymous):

well just do 12*the x values

OpenStudy (anonymous):

so: 12*0=y 12*1=y 12*2=y 12*3=y

OpenStudy (anonymous):

12*3=y is the correct equation to use because x=3

OpenStudy (anonymous):

 that is the table... so 12*0=y 12*1=y 12*2=y 12*3=y 12*4=y?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@tylermcmullen23 im still confused...

OpenStudy (cggurumanjunath):